slope this line 3y =24 - 9x
Answer by 314(4)

in 3y =24 - 9x for X=2, y=(24-9*2)/3 y=(24-18)/3=2 A(2,2) one dot and for x=0 Y=24/3=8 B(0,8) now we can draw the line trough dots A and B
graph(-2,-1,0,1,2,-2,-1,0,1,2,3,4,5,6,7,8, x, (24-9x)/3}A92,2);B(0,8)}}